OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution

OpenStudy (anonymous):

@mjoxprm360

OpenStudy (anonymous):

First subtract 15 from both sides, you get 3/x-8/ = 3 Now divide both sides by 3, you get /x-8/ = 1 Now set x - 8 = 1 and x - 8 = -1 So, x = 9 and x = 7 Seems like the minimum is 7 and maximum is 9.

OpenStudy (anonymous):

@shinebrightlikeadimon

OpenStudy (anonymous):

Okay! What about this one ? Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution

OpenStudy (anonymous):

XD i'm afraid i might get this one wrong again but it's worth a try

OpenStudy (anonymous):

|x - 2| + 10 = 12 |x - 2| = 12-10 |x - 2| = 2 x - 2 = 2 or x - 2 = -2 x = 2+2 or x = -2+2 x = 4 or x = 0

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

|x + 2| + 16 = 14 |x + 2| = 14-16 |x + 2| = -2 Since the absolute value of any number is nonnegative, this means that the last equation is not possible So there are no solutions

OpenStudy (anonymous):

I hope this helped! :)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Anytime :D

OpenStudy (anonymous):

i'm so lost

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Its okay! @mjoxprm360 seems like you know what you are doing!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Your still here? lol